OpenStudy (anonymous):

Use the quadratic formula to find any x-intercepts on the graph of the equation. y=4x^2+8x-6

OpenStudy (anonymous):

this is what I have already got so far before I get stuck

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

a=4 b=8 c=-6

OpenStudy (anonymous):

\[x=\frac{ -8\pm \sqrt{8^{2}-4\times4\times(-6)} }{ 2\times4 }\]

OpenStudy (anonymous):

\[x=\frac{ -8\pm \sqrt{64+96} }{8 }\]

OpenStudy (anonymous):

\[x=\frac{ -8\pm \sqrt{160} }{ 8 }\]

OpenStudy (anonymous):

The I am stuck

OpenStudy (anonymous):

We have to simplify the radical, so what are some factors of 160?

OpenStudy (anonymous):

To make it easier you can use 4*4*10=160 or 4^2*10=160

OpenStudy (anonymous):

1x160 2x80 4x40 5x24 6x20 8x15 10x12

OpenStudy (anonymous):

\[x=\frac{-8\pm 4\sqrt{10} }{ 8}\]

OpenStudy (anonymous):

Do you see how I got that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, One of the factors for 160 is 16*10 We know that the square root of 16 is four, therefore we can remove a four from under the radical, leaving the ten behind.

OpenStudy (anonymous):

Make more sense?

OpenStudy (anonymous):

yes and now I see it because I messed up and was factoring 120 instead of 160

OpenStudy (anonymous):

No problem. You know what to do next?

OpenStudy (anonymous):

it would make your work easier by a factor of two if you started with \[2x^2+4x-3=0\] and then used the quadratic formula

OpenStudy (anonymous):

@satellite73 You wanna help them? I don't want to confuse them by contradicting you? :c

OpenStudy (anonymous):

no you go ahead

OpenStudy (anonymous):

I now simplify by factoring out a common factor of 4 from the numerator and cancelling it out with a factor of 4 in the denominator

OpenStudy (anonymous):

Perfect, so what do you have left?

OpenStudy (anonymous):

\[x=\frac{ -2+\sqrt{10} }{2 }\]

OpenStudy (anonymous):

\[x=\frac{ -2-\sqrt{10} }{ 2 }\]

OpenStudy (anonymous):

Okay so now your x intercepts are \[(\frac{ -2+ \sqrt{10}}{2 },0) and (\frac{ -2- \sqrt{10}}{2 }, 0)\] Unless you're allowed to use a calculator then you can simplify this c:

OpenStudy (anonymous):

ok thanks for helping me.

OpenStudy (anonymous):

No prob c: you already had it

OpenStudy (anonymous):

i got stuck at the factoring but i see where i was messing up at